Operations Report – February 4th

Crew 221 Operations Report 04-02-2022

SOL: 2

Name of person filing report: Luis D. Monge

Non-nominal systems: MDRS Generator is non-operational and still on inspection

Loft water tank connection presents a minor leak

Notes on non-nominal systems: Since Jan 21st, we have a 15 kW rental generator connected to our power system, meanwhile, we are working to troubleshoot the MDRS generator.

A male to male connection was bought in order to replace it in the line, but the leak it’s not in the connection, it’s in the pipe itself.

Generator: off as we are still waiting for it to reach 80%SOC to be turned on

Hours run: 12:00

From what time last night: 8:00pm

To what time this morning: 8:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 89%

Diesel Reading – 22%

Station Propane Reading – 65%

Ethanol Free Gasoline – gallons

Water (loft tank): 32 gallons

Water Meter: 148 units

Water (static tank) – 420 gals

Static to Loft Pump used – yes

Water in GreenHab – 267.4 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: no

Hours: (before EVA) : 128.7

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 134,8

Beginning charge:

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: Notification received on Router replacement, and it will be performed tomorrow as per instructions.

Summary of suits and radios: Radios were used during the EVAs and batteries are malfunctioning and need to be replaced.

Summary of Hab operations: during the afternoon due to high winds and low temperatures the hab temperature dropped from 22C to 18C, the wall heater in the lower deck was turned on and the temperature is back to normal, reading is 20.6C in the thermometer in the wall.

Summary of GreenHab operations: operational, watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators running

Summary of RAM operations: not used

Summary of any observatory issues: no issues

Summary of health and safety issues: no issues

Questions, concerns and requests to Mission Support: a measuring cylinder and two mason jars will be needed for the science experiments, permission to use them is requested.

Copyright © The Mars Society. All rights reserved. | Main Site